OpenStudy (anonymous):

Mr. Mancuso plants tomatoes on 1/3 acre of land, corn on 3/4 acre, and carrots on 1/2 acre. On how many acres of land total did he plant?

OpenStudy (anonymous):

1/3 + 3/4 + 1/2 = 19/12 acre

OpenStudy (anonymous):

simplest form?

OpenStudy (anonymous):

this is it if you want me to approximate it its 19/12= 1.58

OpenStudy (anonymous):

it was wrong

OpenStudy (anonymous):

1 acre and 7/12

OpenStudy (anonymous):

do u mean this way